A 52.8 g ball of copper has a net charge of 2.7 µC. What fraction of the copper's electrons have been removed? (Each copper atom has 29 protons, and copper has an atomic mass of 63.5.)

What is the charge on the sphere contributed by protons? How many total protons are there?
Answer #1

the number of mole for 52.8g copper

52.8/63.5 = 0.8315 mol

the number of copper atoms in 52.8 g copper is

0.8315*6.022x10^23 = 5.007 x10^23

since each atom has 29 proton and electrons, the number of electrons and atoms in 52.8 g copper is

5.007*29*10^23 = 1.452 x10^25

excess number of protons in the ball

2.7 uC/ 1.602x10^-19 = 2.7x10^-6/1.602x10^-19 = 1.685 x10^13

that is the number of e- have been removed

thus the fraction is

1.685x10^13/1.452x10^25

= 1.16 x10^-12

the charge on the sphere contributed by protons is

1.602x10^-19 *number of protons = 1.602x10^-19 *1.452 x10^25 = 2.326 x10^6 C

the total number of protons is 1.452 x10^25